Calculate [H3O+] and [OH−] for each of the following solutions at 25 ∘C given the pH.

Part A) pH= 8.61
Part B) pH= 11.23
Part C) pH= 2.94

Express your answer using two significant figures. Enter your answers numerically separated by a comma.
1
Expert's answer
2020-06-18T14:43:01-0400

Answer.

(a) (first part)

We all know at 250C , pH + pOH = 14, pH =-log[H3O+] and pOH =-log[OH-]

pH = 8.61

so, pOH = (14 -8.61) =5.39

[H3O+] = 10(-8.61) = 2.45 x 10-9 (M)

[OH-] = 10(-5.39) = 4.07 x 10-6 (M)

(b) (second part)

We all know at 250C , pH + pOH = 14, pH =-log[H3O+] and pOH =-log[OH-]

pH = 11.23

so, pOH = (14 -11.23) =2.77

[H3O+] = 10(-11.23) = 5.88 x 10-12 (M)

[OH-] = 10(-2.77) = 1.69 x 10-3 (M)

(c) (third part)

We all know at 250C , pH + pOH = 14, pH =-log[H3O+] and pOH =-log[OH-]

pH = 2.94

so, pOH = (14 -2.94) =11.06

[H3O+] = 10(-2.94) = 1.14 x 10-3 (M)

[OH-] = 10(-11.06) = 8.70 x 10-12 (M)